OpenStudy (anonymous):

Can Someone Help Me Solve : 2x^2-4x-6=0 ?

OpenStudy (lalaly):

\[\large{2(x^2-2x-3)=0}\]\[\large{2(x-3)(x+1)}=0\]

OpenStudy (anonymous):

But My Teacher Doesn't Have That Answer On The Answer Sheet , So That's Why I'm Confused .

OpenStudy (lalaly):

did your teacher ask you to complete the square or factor?

OpenStudy (anonymous):

He Just Said Solve It .

OpenStudy (lalaly):

ok what are the answers u have?

OpenStudy (anonymous):

When I Asked In Class , Another Student Said : 3(2x+3)(2x-3) ?

OpenStudy (lalaly):

no thats wrong, try expanding the brackets you get 12x^2-27 .. and thats not the same as ur equation

OpenStudy (anonymous):

That's What I Said . But The Only Other Choices I Have Too Pick From Is : 2(2x+1)(5x+2) And 5x^2+11x+3 .

OpenStudy (anonymous):

pick none of the above

OpenStudy (lalaly):

are you still tlking about the same question ? because u can solve it in two ways the one i showed u above or u can solve it by copleting the square \[2(x^2-2x-3)=0\]\[2[(x-1)^2-1-3]=0\]\[2(x-1)^2-8=0\]\[(x-1)^2=\frac{8}{2}\]\[(x-1)^2=4\]\[x-1=\pm \sqrt4\]\[x=1+\sqrt4 = 1+2 =3\]and \[x=1-\sqrt4= 1-2= -1\]

OpenStudy (anonymous):

Oh Gosh , I'm SOOO Confused .

OpenStudy (anonymous):

that's just a different way of solving it

OpenStudy (lalaly):

if -1 and 3 are not one of ur choices then ur choices are wrong

OpenStudy (anonymous):

I Have Those Two ( -1 And 3 ) On My Paper , But He Don't Have Them Together . So I Thought I Was Doing It Wrong .